A shopper in a supermarket pushes a cart
with a force of 38 N directed at an angle of
26.4◦ downward from the horizontal.
Find the work done by the shopper as she
moves down a 49.2 m length of aisle.
Answer in units of kJ.
1
Expert's answer
2010-10-05T09:59:21-0400
As you know work is the force multiplies the distance.
In this case we need consider only the horizontal component of force
Fh=F*sin α=38*sin 26.4=16.9
because there no any moving in vertical plane.
Then the work is A=Fh*l=16.9*49.2=831J=0.83kJ

Answer 0.83kJ